# USTC-ICTS/PCFT-25-37

### v1

@@ IV Example 1: the top sector of sunset diagrams
$+\left\{-{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{011}|_{\eta_{1}=0}+{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{101}|_{\eta_{2}=0}\right\},\penalty 10000\ \penalty 10000\ \penalty 10000\$
where $f_{\pm\pm\pm}=\pm m_{1}^{2}\pm m_{2}^{2}\pm m_{3}^{2}-K^{2}$. The unreduced set under these six reduction rules is
${\cal U}=(0,0,0,n_{4},n_{5})\cup(n_{1},n_{2},n_{3},0,0)\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\$(12)
Now we generate new equations from the above six equations by applying differential operators $\partial\over\partial\eta_{i}$}. Among $30$ equations, only $8$ independent new equations are kept according to the step (B), which is the consequence of integrability conditions like ${\partial\over\partial\eta_{k}}\left({\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{k}}\left({\partial\over\partial\eta_{k}}{\partial\over\partial\eta_{i}}\right)=0$. The behavior of new equations depend on the masses in a nontrivial way.
IV.1 The massless case $m_{i}=0,i=1,2,3$
Consider the set $(0,0,0,n_{4},n_{5})$, using $8$ equations we can find two degree-1 type-II equations:
$2\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}+2\eta_{5}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}-(6-3D){\partial\over\partial\eta_{4}}G_{111}$ (13)
$=$ $\left\{{K^{2}(D-2)\over s_{0}}G_{111}+{K^{2}\over s_{0}}\eta_{5}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{K^{2}\over s_{0}}(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G_{111}\right\}+\left\{-{(K^{2})^{2}\eta_{5}\over 2s_{0}^{2}}G_{111}\right\}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$2\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}+2\eta_{4}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}-(6-3D){\partial\over\partial\eta_{5}}G_{111}$ (14)
$=$ $\left\{{K^{2}\eta_{4}\over s_{0}}{\partial\over\partial\eta_{4}}G_{111}+{K^{2}(\eta_{5}+\eta_{4})\over s_{0}}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{K^{2}(D-2)\over s_{0}}G_{111}\right\}+\left\{-{(K^{2})^{2}\eta_{4}\over 2s_{0}^{2}}G_{111}\right\}\penalty 10000\ .\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using above two reduction rules we can reduce any point $(0,0,0,n_{4},n_{5})$ down to $(0,0,0,0,0)$.
Now consider the reduction of the set $(n_{1},n_{2},n_{3},0,0)$, we can form two degree-1 type-III equations:
$\frac{D-4}{2}{\partial\over\partial\eta_{i}}G_{111}-\eta_{i}{\partial^{2}\over\partial\eta_{i}^{2}}G_{111}$ (15)
$=$ $\frac{D-4}{2}{\partial\over\partial\eta_{3}}G_{111}-\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}+\mathcal{R},\ i=1,2\penalty 10000\ \penalty 10000\ \penalty 10000\$
where $\mathcal{R}$ represents lower degree terms. Under the reduction rules produced, the only unreduced set is $(0,0,n_{3},0,0)$. To finish the reduction, we need to generate another equation, which can be achieved by $\partial\over\partial\eta_{3}$} on (13) (or (14)). After simplification we will get
${(D-4)K^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}-{2K^{2}\over s_{0}}\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}=-2\eta_{3}^{2}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
${(5D-16)\eta_{3}}{\partial\over\partial\eta_{3}}G_{111}-{(D-3)(3D-8)}G_{111}.\penalty 10000\ \penalty 10000\ \penalty 10000\$(16)
All nonzero $n_{3}$ can be reduced to $n_{3}=0$ by the reduction rule. Eventually, the only point can not be reduced is $(0,0,0,0,0)$, which is the master integral for the massless case.
IV.2 The massive case
The reduction of the set $(0,0,0,n_{4},n_{5})$ is similar to the massless case, i.e., we can construct equations like (13) and (14). However, to reduce the set $(n_{1},n_{2},n_{3},0,0)$, there are only $5$ independent degree-2 equations with six degree operators ${\partial^{2}\over\partial\eta_{i}^{2}},i=1,2,3$ and ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}},1\leq i<j\leq 3$. Again, after acting $\partial\over\partial\eta_{1}$} on (13), we can form the sixth equation to solve six degree two operators. The final result is that there are four points can not be reduced, i.e., (0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0)(0,0,0,0,0),(1,0,0,0,0),(0,1,0,0,0),(0,0,1,0,0), which give the master integrals for massive sunset.
V Example 2: top sector of two-loop massless double-box
where the last two are ISPs.
There are $10$ fundamental IBP relations. From them, we can solve following $4$ degree-1 operators ${\partial\over\partial\eta_{i}},i=1,3,4,6$ and $6$ degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{j}},i=2,5,7;j=8,9$. The unreduced sets under these $10$ reduction rules are
${\cal U}_{1}$ $=$ $(0,n_{2},0,0,n_{5},0,n_{7},0,0),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9})\penalty 10000\ \penalty 10000\ \penalty 10000\ ,$ (18)
which are similar with the cases of top-sector of sunset. By the step (B), We can generate $8$ independent equations by applying differential operators to $6$ degree-2 equations.
Three of them, i.e., ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{i}}\right)=0,i=2,5,7$, can be applied to produce recurrence relation for $F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}}$, which reduce the set ${\cal U}_{2}$ to point $n_{8}=n_{9}=0$. Considering other $5$ equations for reducing the set ${\cal U}_{1}$, $3$ of them are degree-2 type-I equations for operators ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}},{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}},{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}$. Another two give degree-1 type-III equations
${{D-6}\over 2}{\partial\over\partial\eta_{i}}G-\eta_{i}{\partial^{2}\over\partial\eta_{i}^{2}}G$ (19)
$=$ ${{D-6}\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G+\mathcal{R},\ i=2,7\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Iteratively using reducing rules from these equation, the unreduced set is just $(0,0,0,0,n_{5},0,0,0,0)$. By acting $\partial\over\partial\eta_{5}$} on (19) with $i=2$, and using the reduction rule of ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}$ just found, we can get a recurrence relation for $F_{\vec{n}+2\vec{e}_{5}}$, which reduce any points of $n_{5}\geq 2$. Now the only points can not be reduced is $n_{5}=0,1$, which are two master integrals for top-sector of massless double-box.
VI Example 3: top sector of two-loop massless non-planer
where the last two are ISPs.
There are $10$ fundamental IBP relations. Using them we can solve two degree-1 operators $\partial\over\partial\eta_{1}$}, $\partial\over\partial\eta_{3}$} and $6$ degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{8}},i=2,4,5,6,7$ and ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}$. There are still four unsolved degree-2 operators ${\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{9}},i=4,5,6,7$, which appear in two equations in group form, i.e., $\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{9}}$ and $\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{9}}$. Adding two new integrability conditions, i.e., ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{4}}+{\partial\over\partial\eta_{5}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$ and ${\partial\over\partial\eta_{2}}\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{9}}-\left({\partial\over\partial\eta_{6}}+{\partial\over\partial\eta_{7}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{9}}=0$, we can solve these four operators. Using these reduction rules, now the unreduced sets becomes
${\cal U}_{1}$ $=$ $(0,n_{2},0,n_{4},n_{5},n_{6},n_{7},0,0),$
${\cal U}_{2}$ $=$ $(0,0,0,0,0,0,0,n_{8},n_{9})$ (21)
which becomes the familiar pattern, like the one meet in the top sector of sunset.
Now we do similar procedure. Using the new equations coming from integrability condition ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{8}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{9}}\right)=0,i=2,4,5,6,7$, we can reduce the set ${\cal U}_{2}$ to point $n_{8}=n_{9}=0$. Using the new equations coming from integrability condition ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{i}}{\partial\over\partial\eta_{k}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{k}}\right)=0,i<j,i,j=2,4,5,6,7;k=8,9$, we can solve degree two operators ${\partial\over\partial\eta_{j}}{\partial\over\partial\eta_{i}}$ and four degree one operator ${\partial\over\partial\eta_{i}},i=2,5,6,7$. Finally by generating a new degree two equation for $\partial^{2}\over\partial\eta_{4}^{2}$}, we are able to reduce the set ${\cal U}_{1}$ to two points $n_{2}=n_{5}=n_{6}=n_{7}=0$, $n_{4}=0,1$, i.e., two master integrals. For more details, see the supplemental material.
VII Conclusion and Outlook
In this work, we introduced a novel systematic approach to derive complete symbolic reduction rules for multi-loop Feynman integrals with generating functions. By the DEs for these generating functions, we obtain recurrence relations that enable the efficient reduction without the exponential proliferation of IBP identities in traditional methods. Through explicity examples, including the sunset diagrams (both massive and massless), two-loop double-box, and two-loop non-planar integrals, we demonstrate the algorithm’s effectiveness in complex cases with high-degree numerators and high-power propagators. This method addresses key bottlenecks in perturbative QFT computations.
With the proof of concept in the note, the most important thing is to automate the aforementioned algorithm. In a longer companion paper, more details of the algorithm will be presented. The combination of the idea in the paper and other reduction methods is also worth pursuing.
Acknowledgements.
Supplemental Material
In the supplemental material, we provide some details of the examples in this paper.
Appendix A The topsector of sunset
Using six DEs coming from six fundamental IBP relations, we can solve six degree two operators as
@@ Appendix A The topsector of sunset
where since we focus on the top sector only, contributions from subsectors have been neglected. Also, we have written into the form with degree separated. They give six good reduction rules. Using them, we can reduce any point except the following two sets ${\cal U}_{1}=(n_{1},n_{2},n_{3},0,0)$ and ${\cal U}_{2}=(0,0,0,n_{4},n_{5})$. To reduce the two remaining sets, we need to generate new equations. Acting on $\partial\over\partial\eta_{i}$} on above six equations, we will get $30$ degree $3$ equations. However, by using descendant reduction rules of above six reduction rules, one can check that we get $9$ degree $2$ equations, where one of them is
@@
$\left.+(4-2D){\partial\over\partial\eta_{5}}G_{111}\right\}$
$+\left\{+{2(m_{2}^{2}-m_{3}^{2})\eta_{1}+2m_{1}^{2}(\eta_{2}-\eta_{3})\over s_{0}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-{m_{2}^{2}(\eta_{1}+\eta_{2}-\eta_{3})+f_{++-}\eta_{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$+{-f_{+-+}\eta_{3}-m_{3}^{2}(\eta_{1}-\eta_{2}+\eta_{3})\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}$
$+{2K^{2}-f_{-+-}\over 2s_{0}}\eta_{4}{\partial\over\partial\eta_{4}}G_{111}$
$\left.+{2(K^{2}+m_{1}^{2})\eta_{4}-f_{++-}\eta_{5}\over 2s_{0}}{\partial\over\partial\eta_{5}}G_{111}\right$.
$\left.+{(D-2)K^{2}-(D-3)(m_{2}^{2}-m_{3}^{2})\over s_{0}}G_{111}\right\}$
$+\left\{+2(\eta_{2}-\eta_{3})\eta_{1}{\partial\over\partial\eta_{1}}G_{111}+(\eta_{1}+\eta_{2}-\eta_{3})\eta_{2}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.-(\eta_{1}-\eta_{2}+\eta_{3})\eta_{3}{\partial\over\partial\eta_{3}}G_{111}+{(\eta_{1}-\eta_{2}+\eta_{3})\over 2}\eta_{4}{\partial\over\partial\eta_{4}}G_{111}\right$.
$+(\eta_{1}\eta_{4}-{1\over 2}\eta_{5}(\eta_{1}+\eta_{2}-\eta_{3})){\partial\over\partial\eta_{5}}G_{111}$
$\left.+({K^{2}f_{--+}\over 2s_{0}^{2}}\eta_{4}+{K^{2}m_{2}^{2}\over s_{0}^{2}}\eta_{5}-(D-3)(\eta_{2}-\eta_{3}))G_{111}\right\}$
$+\left\{+({K^{2}\over s_{0}}\eta_{2}\eta_{5}-{K^{2}\over 2s_{0}}\eta_{4}(\eta_{1}+\eta_{2}-\eta_{3}))G_{111}\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$
@@
These equations are, in fact, the consequence of integrability conditions ${\partial\over d\eta_{i}},{\partial\over\partial\eta_{j}}]=0$. Above $9$ new equations can be divided into three groups. The first group is given by ${\partial\over\partial\eta_{4}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{5}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)=0$ with $i=1,2,3$. We will denote the corresponding equations as (I-i) ((28) is (I-1)). The second group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (II-ij). The third group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (III-ij). These $9$ equations are not independent since  ${\text{ (II-12)}}+{\text{ (III-12)}}-{\text{ (II-23)}}-{\text{ (III-13)}}=0$. Also the behavior of new equations depend on the masses in a nontrivial way.
The result (28) is presented according to the degree of operators. For later discussions, only the degree one and two operators are important, so we will neglect the other parts. Similarly, the relevant part of (I-2) is
$0$ $=$ $\left\{-\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}-2\eta_{5}{\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}G_{111}\right$. (29)
$\left.+{2m_{2}^{2}(m_{1}^{2}-m_{3}^{2})\over s_{0}^{2}}{\partial\over\partial\eta_{2}}G_{111}+{m_{1}^{2}f_{++-}\over s_{0}^{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-{m_{3}^{2}f_{-++}\over s_{0}^{2}}{\partial\over\partial\eta_{3}}G_{111}+(4-2D){\partial\over\partial\eta_{4}}G_{111}\right$.
$\left.+(2-D){\partial\over\partial\eta_{5}}G_{111}\right\}+.$$penalty 10000\ \penalty 10000\$
and the relevant part of (I-3) is
$0$ $=$ $\left\{-\eta_{4}{\partial^{2}\over\partial\eta_{4}^{2}}G_{111}+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G_{111}+{m_{1}^{2}f_{+-+}\over s_{0}^{2}}{\partial\over\partial\eta_{1}}G_{111}\right$. (30)
$-{m_{2}^{2}f_{-++}\over s_{0}^{2}}{\partial\over\partial\eta_{2}}G_{111}+{2m_{3}^{2}(m_{1}^{2}-m_{2}^{2})\over s_{0}^{2}}{\partial\over\partial\eta_{3}}G_{111}$
$\left.-(D-2){\partial\over\partial\eta_{4}}G_{111}+(D-2){\partial\over\partial\eta_{5}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The special point of the above three equations is that the highest operators have degree one. Groups II and III are degree two equations. The (II-12) is
${2m_{2}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}+{f_{+++}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (31)
$+{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}$
$=$ $\left\{-2\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G_{111}-(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$\left.+(D-3){\partial\over\partial\eta_{1}}G_{111}+(2D-7){\partial\over\partial\eta_{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (II-13) is
${2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}+{2m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (32)
$+{f_{+++}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$
$=$ $\left\{(D-3){\partial\over\partial\eta_{1}}G_{111}+(2D-7){\partial\over\partial\eta_{3}}G_{111}-2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.
$\left.-2\eta_{2}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}-2\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.-(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (II-23) is
@@
The (III-12) is
${-2m_{1}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}-{f_{+++}\over s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$ (34)
$-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}-{2m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}$
$=$ $\left\{2\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}+(\eta_{1}+\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}+2\eta_{3}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{2}}G_{111}\right$.
@@
$\left.+(7-2D){\partial\over\partial\eta_{1}}G_{111}-(D-3){\partial\over\partial\eta_{2}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
The (III-13) is
$-{2m_{1}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}+{2m_{3}^{2}\over s_{0}}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}$ (35)
$=$ $\left\{(D-4){\partial\over\partial\eta_{3}}G_{111}-(D-4){\partial\over\partial\eta_{1}}G_{111}-2\eta_{3}{\partial^{2}\over\partial\eta_{3}^{2}}G_{111}\right$.

### v2

@@ Appendix A The topsector of sunset
$-\left\{-{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{011}|_{\eta_{1}=0}+{1\over s_{0}}{\partial\over\partial\eta_{3}}G_{101}|_{\eta_{2}=0}\right\},\penalty 10000\ \penalty 10000\ \penalty 10000\$(29)
@@
where since we focus on the top sector only, contributions from subsectors have been neglected. Also, we have written into the form with degree separated to demonstrate our idea.
Using them, we can reduce any point except the following two sets ${\cal U}_{1}=(n_{1},n_{2},n_{3},0,0)$ and ${\cal U}_{2}=(0,0,0,n_{4},n_{5})$. Since there are still infinite number of irreducible points, we need go to the next round of computation.
The second round of computation: Acting on $\partial\over\partial\eta_{i}$} on above six equations, we will get $30$ degree-3 equations. However, among them, only $9$ provides nontrivial new equations. To see it, let us consider the equation produced by $\partial\over\partial\eta_{4}$} on the reduction rule ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{1}}=$.... We get a degree-3 operator ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{1}}$, which is the descendant of another operator ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{1}}$ and can be simplified using the reduction rule (32) with the action $\partial\over\partial\eta_{5}$}. After done it, the equation becomes degree-2 with following operators ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{3}}$, ${\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{1}}$, ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{1}}$, ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}$ and ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{3}}$. All of them can be simplified and we reach a new degree-1 equation
@@
$\left.+(4-2D){\partial\over\partial\eta_{5}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
where we have kept only the leading degree part.
As can see from this example, only after this simplification, the dust coming from the known reduction rules will be cleaned away and new unsolved operators will appear manifestly as the leading degree operators. Solving these new appeared operators we will get new reduction rules. Doing it round by round, eventually we will find the complete reduction rules.
@@
These equations are, in fact, the consequence of integrability conditions ${\partial\over d\eta_{i}},{\partial\over\partial\eta_{j}}]=0$. Above $9$ new equations can be divided into three groups. The first group is given by ${\partial\over\partial\eta_{4}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{5}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)=0$ with $i=1,2,3$. We will denote the corresponding equations as (I-i) ((36) is (I-1)). The second group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{4}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (II-ij). The third group is given by ${\partial\over\partial\eta_{j}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{i}}\left({\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{j}}\right)=0$ with $1\leq i<j\leq 3$, which will be denoted as (III-ij). These $9$ equations are not independent since  ${\text{ (II-12)}}+{\text{ (III-12)}}-{\text{ (II-23)}}-{\text{ (III-13)}}=0$.
Another important feature is that the degree of equations depend on the masses. For example, the (II-23) is
@@
which will be degree-1 for massless case. This difference will lead to different master integrals.
Solving module of second round for massless case: For $m_{i}=0,i=1,2,3$, we have $3$ degree-2 equations with coefficient matrix as
$\left(\begin{array}[]{c| c |c |l}[1,2]&[1,3]&[2,3]&\\ \hline\cr{K^{2}\over s_{0}}&&&{\rm(II-12)}\\ \hline\cr&{K^{2}\over s_{0}}&&{\rm(II-13)}\\ \hline\cr&&-{K^{2}\over s_{0}}&{\rm(III-23)}\end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(42)
and $5$ degree-1 equations with coefficient matrix is
@@
$\left(\begin{array}[]{c| c |c |c| c| l}{\partial\over\partial\eta_{1}}&{\partial\over\partial\eta_{2}}&{\partial\over\partial\eta_{3}}&{\partial\over\partial\eta_{4}}&{\partial\over\partial\eta_{5}}&\\ \hline\cr&&&-\widehat{O}_{0;4,1}&\widehat{O}_{0;5,2}&{\rm(I-1)}\\ \hline\cr&&&\widehat{O}_{0;4,2}&-\widehat{O}_{0;5,1}&{\rm(I-2)}\\ \hline\cr&&&-\widehat{O}_{0;4,1}&\widehat{O}_{0;5,1}&{\rm(I-3)}\\ \hline\cr&\widehat{O}_{0;2,1}&-\widehat{O}_{0;3,1}&&&{\rm(II-23)}\\ \hline\cr-\widehat{O}_{0;1,1}&&\widehat{O}_{0;3,1}&&&{\rm(III-13)}\\ \end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(49)
with
$\widehat{O}_{0;a,1}$ $=$ $(D-2)+\eta_{a}{\partial\over\partial\eta_{a}},\penalty 10000\ \penalty 10000\ ,a=4,5$
$\widehat{O}_{0;4,2}$ $=$ $4-2D-\eta_{4}{\partial\over\partial\eta_{4}}-2\eta_{5}{\partial\over\partial\eta_{5}}$
$\widehat{O}_{0;5,2}$ $=$ $4-2D-\eta_{5}{\partial\over\partial\eta_{5}}-2\eta_{4}{\partial\over\partial\eta_{4}}$
$\widehat{O}_{0;i,1}$ $=$ $(D-4)-2\eta_{i}{\partial\over\partial\eta_{i}},\penalty 10000\ i=1,2,3\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(50)
Doing Gauss elimination for the second and third row gives